# On the smoothness of solutions of fully nonlinear second order equations in the plane

### v1

@@ 1. Introduction
  • If $n\geq 5$ there are counterexamples to the $C^{1,1}$ regularity [NTV14], and the best regularity is $C^{1,\alpha}$ by fundamental results proved by L. Caffarelli [Caf89, CC95] and N. Trudinger [Tru88]. The cases $n=3,4$ remain at this stage the big open problems in the theory: solutions are known to be $C^{1,\alpha}$ under the sole uniform ellipticity, but we do not know whether they satisfy better regularity, cf. [NTV14].
  • If $F$ is convex or concave (i.e. Bellman equations), solutions are known to be $C^{2,\alpha}$, $\alpha\in(0,1)$ small, by a result of L.C. Evans [Eva82] and N.V. Krylov [Kry82]. This was earlier proved through the theory of variational inequalities by H. Brézis and L.C. Evans [BE79] for the maximum of two operators. If $F$ is only continuous the best regularity of Bellman equations is one of the major open problems. When $n=2$ this was solved in the case $F(D^{2}u)=\min\{L_{1}u,L_{2}u\}$ with $L_{1},L_{2}$ having constant coefficients, and solutions of these special Bellman PDEs are $C^{2,1}$, as proved by L. Caffarelli-D. De Silva-O. Savin [CDSS18]. Higher smoothness than $C^{2,\alpha}$ ($\alpha$ implicit) remains open when $n\geq 3$.
@@
H. Brézis [Bré78, Comments and open questions- (2)] and H. Brézis-L.C. Evans [BE79, Remark 4.2] asked whether one can expect better regularity than $C^{2,\alpha}$ for $\alpha$ small or $W^{3,p}$ for some $p$. The Hölder exponent is implicit since one usually applies the De Giorgi-Nash-Moser or Krylov-Safonov theory to the equation satisfied by the second derivatives. $C^{2,\alpha}$ smoothness is false when $n\geq 5$ by the counterexamples in [NTV14], but it remains an interesting question in the lower dimensional cases $n\leq 4$, starting from the plane $\mathbb{R}^{2}$, and for convex equations. We are only aware of a result by Q. Huang [Hua02] showing that if $F$ is uniformly elliptic, then $u\in C^{2,\alpha}$ for any $\alpha\in(0,1)$ and any $n\geq 2$, with estimates depending on the first derivatives of $F$, whenever the equation has a polynomial Liouville property. In the special case $F(D^{2}u)=\min\{L_{1}u,L_{2}u\}$ an optimal $C^{2,1}$ estimate is given in [CDSS18].
As far as $W^{3,p}$ estimates are concerned, these are known for solutions of (1) in any dimension for an implicit $p<1$ depending on $n$ and $\lambda/\Lambda$, as a consequence of [Le20, Theorem 1.4] or [ASS12]. It is enough to apply the $W^{2,\varepsilon}$-estimate of [Eva85, CC95, Lin86] to the equation solved by the directional derivative $\partial_{e}u$, for which estimates of the decay of the integrability are available [ASS12, Le20, Moo19, NT23]. The authors in [ASS12] observed that $\varepsilon=\frac{2\lambda}{\Lambda+\lambda}$ is the universal upper bound in the two-dimensional case for solutions of $\mathcal{M}^{+}_{\lambda,\Lambda}(D^{2}u)\geq 0$. However, this exponent can be improved when $\lambda=\Lambda$, where $\varepsilon=1$.
Our paper gives some results in the first direction, proving some $C^{2,\alpha}$ estimates and providing an explicit Hölder exponent depending on $\lambda/\Lambda$. We use the linear theory of equations in divergence and nondivergence form and exploit some known regularity results that depend only on the bounds of the coefficients, but not on their regularity properties. To our knowledge these results have not appeared in print anywhere. We first prove that, in the general case, $C^{2}$ solutions of (1) when $n=2$ are more regular and satisfy
@@
with $\bar{\alpha}$ depending on $\lambda$ and $\Lambda$. These results use that the second derivatives solve a suitable equation in divergence form. We then improve this bound via the theory of equations in nondivergence form, proving
@@
It is worth mentioning that there is a counterpart of these results for nonlinear equations in divegence form, whose main model is the $p$-Laplacian. An explicit Hölder exponent, depending on $p$, for the regularity of the gradient was proved in [IM89]: solutions of $p$-harmonic functions in the plane are $C^{1,\alpha}$ with
@@
Outline. In Section 2 we recall the notion of uniform ellipticity for (1). Section 3 provides a short survey of regularity properties of solutions to equations in divergence and nondivergence form in the plane. Section 4 exploits the result for equations in divergence form to prove $C^{2,\alpha}$ and $W^{3,p}$ estimates for fully nonlinear equations, while Section 5 uses the nonvariational regularity theory in the plane to prove $C^{2,\alpha}$ estimates for a different value of $\alpha$.
@@ 2. Preliminaries on fully nonlinear equations
This holds even if $F$ is not $C^{1}$, since the uniform ellipticity implies the Lipschitz continuity in the matrix entry: in this case the derivatives must be understood almost everywhere. If $F$ is uniformly elliptic according to the previous definition, then the linearization of $F$ is uniformly elliptic with the same ellipticity constants, namely for all $M$
@@ 3.1. Divergence form equations
This result can be improved for equations posed on $\mathbb{R}^{2}$. When $A$ is symmetric, an explicit exponent $\alpha=\frac{\lambda}{2\Lambda}$ was found by C. Morrey [Mor43], while [Wid69] provides the exponent $\alpha=\frac{1}{8}\sqrt{\frac{\lambda}{\Lambda}}$. The sharp result is contained in the following
@@
We also have the following generalization valid for certain non-symmetric matrices
@@ 4. $C^{2,\alpha}$ estimates via the variational regularity theory
First, note that $\mathrm{Tr}(\tilde{A}^{s})>0$ and $\mathrm{det}(\tilde{A}^{s})=\frac{\mathrm{det}(A)}{a_{22}^{2}}>0$, so it is positive definite. Moreover, $\tilde{A}^{s}$ satisfies
@@
we conclude that $w\in C^{0,\bar{\alpha}}(B_{\frac{1}{2}})$ by [Tre74, Tre85] with
@@
One can also reach an explicit Hölder exponent for divergence form equations via the Widman hole-filling technique, by taking care of the constants in the estimates.
@@
which gives the statement after sending $r\to+\infty$, since $Dw=0$ and $w$ is constant in $\mathbb{R}^{2}$. If $F\in C^{1}$, this implies by a result of [Hua02] that $u\in C^{2,\alpha}$ for any $\alpha<1$, with estimates depending on $F$ and the modulus of continuity of $DF$. As pointed out in [Moo24], we also note that the previous argument provides a Liouville property for equations in nondivergence form: $C^{2}$ solutions of $\mathrm{Tr}(A(x)D^{2}u)=0$ with bounded gradient in the plane are linear functions. The previous computations give a new proof of this result.
Remark 4.4.
The Hölder regularity exponent can be improved if we consider equations of the form $f(\Delta u)=g(x)$ with $g$ regular enough (e.g. $\Delta g\in L^{q}$, $q$ sufficiently large) and $\lambda\leq f^{\prime}(\zeta)\leq\Lambda$. In fact, $z=\Delta u$ solves
$\sum_{i=1}^{2}\partial_{i}(f^{\prime}(\Delta u)\partial_{i}z)=\Delta g\in L^{q}.$
By the results for nonhomogeneous equations in divergence form with symmetric matrices one can prove that $u\in C^{0,\alpha}$ with $\alpha=\sqrt{\frac{\lambda}{\Lambda}}$ or $\alpha=\frac{4}{\pi}\arctan\left(\sqrt{\frac{\lambda}{\Lambda}}\right)$, see e.g. [Alv08, PS72, Tre74]. One can also prove $W^{3,p}$ estimates for such equations via the results in [LN97]. One can also proceed in a similar manner for the general higher dimensional case: in this setting there are known estimates of the Harnack constant and the Hölder exponent found in [BG72].
@@ 5. $C^{2,\alpha}$ estimates via the nonvariational regularity theory
We prove now better $C^{2,\alpha}$ estimates using the theory of nondivergence form equations in the plane.
@@
However, when $\frac{\lambda}{\Lambda}\to 1$ we have $\hat{\alpha}\to\frac{\sqrt{33}-3}{2}<1$, while $\tilde{\alpha}\to 1$.

### v2

@@ 1. Introduction
  • If $n\geq 5$ there are counterexamples to the $C^{1,1}$ regularity [NTV14], and the best regularity is $C^{1,\alpha}$ by fundamental results proved by L. Caffarelli [Caf89, CC95] and N. Trudinger [Tru88]. The cases $n=3,4$ remain at this stage the main open problems in the theory: solutions are known to be $C^{1,\alpha}$ under the sole uniform ellipticity, but we do not know whether they satisfy better regularity, cf. [NTV14].
  • If $F$ is uniformly elliptic and either concave or convex (i.e. Bellman equations), solutions are known to be $C^{2,\alpha}$, $\alpha\in(0,1)$ small, by a result of L.C. Evans [Eva82] and N.V. Krylov [Kry82]. This was earlier proved through the theory of variational inequalities by H. Brézis and L.C. Evans [BE79] for the maximum of two operators. If $F$ is only continuous the maximal regularity of Bellman equations is one of the major open problems. When $n=2$ this was solved in the case $F(D^{2}u)=\min\{L_{1}u,L_{2}u\}$ with $L_{1},L_{2}$ having constant coefficients, and solutions of these special Bellman PDEs are $C^{2,1}$, as proved by L. Caffarelli-D. De Silva-O. Savin [CDSS18]. Higher smoothness than $C^{2,\alpha}$ ($\alpha$ implicit) remains unsettled when $n\geq 3$.
@@
H. Brézis [Bré78, Comments and open questions- (2)] and H. Brézis-L.C. Evans [BE79, Remark 4.2] asked whether one can expect better regularity than $C^{2,\alpha}$ for $\alpha$ small and $W^{3,p}$ for some $p$. The Hölder exponent is implicit since one usually applies the De Giorgi-Nash-Moser or Krylov-Safonov theory to the equation satisfied by the second derivatives. $C^{2,\alpha}$ smoothness is false when $n\geq 5$ by the counterexamples in [NTV14], but it remains an interesting question in the lower dimensional cases $n\leq 4$, starting from the plane $\mathbb{R}^{2}$, and for convex equations. We are only aware of a result by Q. Huang [Hua02] showing that if $F$ is uniformly elliptic, then $u\in C^{2,\alpha}$ for any $\alpha\in(0,1)$ and any $n\geq 2$, with estimates depending on the first derivatives of $F$, whenever the equation has a polynomial Liouville property. In the special case $F(D^{2}u)=\min\{L_{1}u,L_{2}u\}$ an optimal $C^{2,1}$ estimate is given in [CDSS18].
As far as $W^{3,p}$ estimates are concerned, these are known for solutions of (1) in any dimension for an implicit $p<1$ depending on $n$ and $\lambda/\Lambda$, as a consequence of [Le20, Theorem 1.4] or [ASS12]. It is enough to apply the $W^{2,\varepsilon}$-estimate of [Eva85, CC95, Lin86] to the equation solved by the directional derivative $\partial_{e}u$, for which estimates of the decay of the integrability are available [ASS12, Le20, Moo19, NT23]. The authors in [ASS12] observed that $\varepsilon=\frac{2\lambda}{\Lambda+\lambda}$ is the universal upper bound in the two-dimensional case for solutions of $\mathcal{M}^{+}_{\lambda,\Lambda}(D^{2}u)\geq 0$. However, this exponent can be improved when $\lambda=\Lambda$, where $\varepsilon=1$. The recent manuscript [NT25] addressed lower bounds for the exponent $\varepsilon$, proving in particular the new estimate $\varepsilon>\frac{1.629\lambda}{\lambda+\Lambda}$.
Our paper gives some quantitative results in the first direction, proving some $C^{2,\alpha}$ estimates and providing an explicit lower bound for the Hölder exponent depending on $\lambda/\Lambda$. This can be considered as a counterpart in $C^{2,\alpha}$ spaces of the results recently obtained in [NT25] for $W^{2,\varepsilon}$ estimates. We use the linear theory of equations in divergence and nondivergence form and exploit some known regularity properties of their solutions that depend only on the bounds of the coefficients, but not on their regularity properties. To our knowledge these results have not appeared in print anywhere. We first prove that, in the general case, $C^{2}$ solutions of (1) when $n=2$ and $F$ is uniformly elliptic are more regular and satisfy
@@
with $\bar{\alpha}$ explicit and depending on $\lambda$ and $\Lambda$. These results use that the second derivatives solve a suitable equation in divergence form, see Section 3. We then improve this bound via the theory of equations in nondivergence form, proving
@@
It is worth mentioning that there is a counterpart of this theory for nonlinear equations in divegence form, whose main model is the $p$-Laplacian. An explicit Hölder exponent, depending on $p$, for the regularity of the gradient of solutions of the $p$-Laplace equation was proved in [IM89]: solutions of $p$-harmonic functions in the plane are $C^{1,\alpha}$ with
@@
Outline. In Section 2 we recall the notion of uniform ellipticity for (1). Section 3 provides a short survey of regularity properties of solutions to equations in divergence and nondivergence form in the plane. Section 4 exploits the result for equations in divergence form to prove $C^{2,\alpha}$ and $W^{3,p}$ estimates for fully nonlinear equations, while the main result in Section 5 uses the nonvariational regularity theory in the plane to prove $C^{2,\alpha}$ estimates for a different value of $\alpha$.
@@ 2. Preliminaries on fully nonlinear equations
This holds even if $F$ is not $C^{1}$, since the uniform ellipticity implies the Lipschitz regularity in the matrix entry: in this case the derivatives must be understood almost everywhere. If $F$ is uniformly elliptic according to the previous definition, then the linearization of $F$ is uniformly elliptic with the same ellipticity constants, namely for all $M$
@@ 3.1. Divergence form equations
This result can be made quantitative for equations posed on $\mathbb{R}^{2}$. When $A$ is symmetric, an explicit exponent $\alpha=\frac{\lambda}{2\Lambda}$ was found by C. Morrey [Mor43], while [Wid69] provided the exponent $\alpha=\frac{1}{8}\sqrt{\frac{\lambda}{\Lambda}}$. The sharp result is contained in the following:
@@
We also have the following generalization valid for certain non-symmetric matrices:
@@ 4. $C^{2,\alpha}$ estimates via the variational regularity theory
First, note that $\mathrm{Tr}(\tilde{A}^{s})>0$ and $\mathrm{det}(\tilde{A}^{s})=\frac{\mathrm{det}(A)}{a_{22}^{2}}>0$, therefore it is positive definite. Moreover, $\tilde{A}^{s}$ satisfies
@@
we conclude that $w\in C^{0,\bar{\alpha}}(B_{\frac{1}{2}})$ by Theorem 3.3, cf. [Tre74, Tre85], with
@@
One can also reach an explicit Hölder exponent, depending on $\lambda/\Lambda$, for divergence form equations via the Widman hole-filling technique [Moo22, Section 3] by taking care of the constants in the estimates.
@@
which gives the statement after sending $r\to+\infty$, since $Dw=0$ and $w$ is constant in $\mathbb{R}^{2}$. If $F\in C^{1}$, this implies by a result of [Hua02] that $u\in C^{2,\alpha}$ for any $\alpha<1$, with estimates depending on $F$ and the modulus of continuity of $DF$. We also note that the previous argument provides a new proof of the Liouville property for homogeneous equations in nondivergence form with bounded and measurable coefficients: $C^{2}$ solutions of $\mathrm{Tr}(A(x)D^{2}u)=0$ with bounded gradient in the plane are linear functions, cf. [Moo24]. In fact, it is enough to divide the equation by $a_{22}(x)\geq\lambda$ and then differentiate with respect to $x_{1}$ to find that $w=\partial_{1}u$ solves $\mathrm{div}(\tilde{A}(x)Dw)=0$ in $\mathbb{R}^{2}$.
@@ 5. $C^{2,\alpha}$ estimates via the nonvariational regularity theory
We prove now better $C^{2,\alpha}$ quantitative estimates using the theory of nondivergence form equations in the plane.
@@
However, when $\frac{\lambda}{\Lambda}\to 1$ we have $\hat{\alpha}\to\frac{\sqrt{33}-3}{4}<1$, while $\tilde{\alpha}\to 1$.